OpenStudy (anonymous):

check my math please differentiate f(x)=((x)/x^3+1)^6

OpenStudy (anonymous):

i got this

OpenStudy (anonymous):

Figured it out?

OpenStudy (anonymous):

\[6((x)/(x^3+1))^5(-2x^3+1)/(x^3+1)^2\]

OpenStudy (anonymous):

i just want u to check my answer please :D

OpenStudy (anonymous):

otherwise correct me pls :D

OpenStudy (anonymous):

is that the guy from digimon :D?

OpenStudy (freckles):

it looks fabulous but you could also use log differentiation

OpenStudy (anonymous):

they havn't teach me that yet :D

OpenStudy (anonymous):

finish what you are typing i got another question :)

OpenStudy (anonymous):

for sin(x+y)=2x-2y if i differentiate this i would get cos(x+y)+1+y' = 2-2y' right ?

OpenStudy (anonymous):

then it says find the equation of the tangent line at the point (pi,pi). and write in slope intercept form

OpenStudy (anonymous):

should i substitute cos(pi+pi)+1+y' = 2-2y' and solve for y'?

OpenStudy (freckles):

\[y=(\frac{x}{x^3+1})^6 \\ \ln(y)=6 \ln(\frac{x}{x^3+1}) \\ \ln(y)=6(\ln(x)-\ln(x^3+1)) \\ \ln(y) =6\ln(x)-6\ln(x^3+1) \] now apply that d/dx ln(u)=u'/u so we have \[\frac{y'}{y}=6 \frac{1}{x}-6 \frac{3x^2}{x^3+1} \\ \frac{y'}{y}=\frac{6}{x}-\frac{18x^2}{x^3+1} \\ \] You could multiply y on both sides and call it quits or multiply y on both sides then replace y with what you had in the first line we could also do a little more work like combining fractions first... \[\frac{y'}{y}=\frac{6(x^3+1)-18x^2(x)}{x(x^3+1)} \\ \frac{y'}{y}=\frac{-12x^3+6}{x(x^3+1)} \\ \text{ multiply y on both sides } y' =y \frac{-12x^3+6}{x(x^3+1)} \\ \text{ recall } y=(\frac{x}{x^3+1})^6 \\ y'=(\frac{x}{x^3+1})^6 \cdot \frac{-12x^3+6}{x(x^3+1)} \\ y'=6 \frac{x^5}{(x^3+1)^6} \frac{-2x^3+1}{x^3+1} \\ y'=6(\frac{x}{x^3+1})^5\frac{-2x^3+1}{(x^3+1)^2}\] just trying to show you get the exact same thing

OpenStudy (freckles):

now back to your other question

OpenStudy (freckles):

it looks like you tried to use chain rule but ummm...

OpenStudy (freckles):

kinda not entirely but yeah kinda failed just a smidgen

OpenStudy (freckles):

copying and pasting: "for sin(x+y)=2x-2y if i differentiate this i would get cos(x+y)+1+y' = 2-2y' right ?" you should put a multiplication operator between cos(x+y) and (1+y')

OpenStudy (freckles):

that is chain rule does derivative of outside +derivative of inside it says derivative of outside*derivative of inside

OpenStudy (freckles):

that is you should have cos(x+y)*(1+y')=2-2y'

OpenStudy (anonymous):

i see

OpenStudy (freckles):

and yes you can replace x with pi and y with pi and solve for y'

OpenStudy (freckles):

and y' at (pi,pi) will be the slope of your line

OpenStudy (anonymous):

cool thanks :D